OpenStudy (anonymous):

1/4(x-12)+3/4=x/3 Solve for X.

OpenStudy (anonymous):

multiply by 12 as a start to clear the fractions

OpenStudy (anonymous):

x12 to every fraction?

OpenStudy (anonymous):

\[\frac{1}{4}(x-2)+\frac{3}{4}=\frac{x}{3}\] \[12\left(\frac{1}{4}(x-2)+\frac{3}{4}\right)=\frac{x}{3}\times 12\] \[12\times \frac{1}{4}(x-2)+12\times \frac{3}{4}=4x\] \[3(x-2)+9=4x\]and it should be good from there

OpenStudy (anonymous):

the point is to cancel all the denominators, which is why you choose 12 as it is the least common multiple of 3 and 4

OpenStudy (anonymous):

x/4-12/4+3/4=x/3 x-12+3/4=x/3 x-9/4=x/3 3(x-9)=4x 3x-27=4x x=-27